Circular No. 6251

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

51 PEGASI M. Mayor and D. Queloz, Geneva Observatory, have reported the dis-

covery of a Jupiter-mass object in orbit around the solar-type star 51 Peg. The announcement was made in Florence on Oct. 6 at the Ninth “Cambridge” Workshop on “Cool Stars, Stellar Systems, and the Sun”. The claim is based on 18 months of precise Doppler measurements made with the ELODIE spectrograph of the Observatoire de Haute-Provence. The parameters of the orbital motion are as follows: P = 4.2293 ± 0.0011 days, e = 0 (assumed), K = 0.059 ± 0.003 km/s, T_{0} = 2449797.773 ± 0.036. The minimum mass of the companion is 0.47 ± 0.02 Jupiter mass. Alternative explanations for the radial-velocity variation (pulsation or spot rotation) seem to be ruled out by the absence of any significant corresponding photometric variation.

Following the Oct. 6 announcement, confirmation of the 4.2-day radialvelocity variation was obtained in mid-October by G. Marcy and P. Butler (San Francisco State University, University of California at Berkeley) at the Lick Observatory, as well as by a joint team from the Harvard-Smithsonian Center for Astrophysics (R. Noyes, S. Korzennik, M. Krockenberger and P. Nisenson), the High Altitude Observatory (T. Brown, T. Kennelly and C. Rowland) and Pennsylvania State University (S. Horner).

G. Burki, M. Burnet and M. Kuenzli, Geneva Observatory and Lausanne University, communicate: “Intensive photometric monitoring of 51 Peg has been carried out at the European Southern Observatory. There is no evidence for eclipses in the system. The rms of the V magnitude (on 17 nights) is 0.037, two comparison stars being used. A 4.2-day photometric variability larger than 0.002 mag can be ruled out.”

COMET C/1995 Q1 (BRADFIELD) Total magnitude estimates, visual unless otherwise stated (B = binoc-

ulars): Sept. 26.95 UT, 8 (Zhu Xinmin, Xinjiang Province, China; independent discovery); 30.20, 8.1 (J. Carvajal, Madrid, Spain, 0.44-m reflector); Oct. 1.15, 7.6, (R. Haver, Frasso Sabino, Italy, 15 × 80 B); 5.18, 8.3 (J. D. Shanklin, Cambridge, England, 14 × 100 B); 7.44, 8.2 (G. W. Kronk, Troy, IL, 20×80 B); 14.16, 9.2 (H. Mikuž, Ljubljana, Slovenia, 0.19-m reflector + CCD + V filter); 19.24, 9.3: (A. Pereira, Cabo da Roca, Portugal, 0.15-m reflector); 21.50, 8.3 (C. S. Morris, Pine Mountain Club, CA, 20 × 80 B); 24.11, 9.2 (G. A. Milani, Mt. Tomba, Italy, 20 × 80 B).

1995 October 25 Brian G. Marsden